OpenStudy (anonymous):

g(s)=(s+4)/((s^2)+1s+14) find g'(s). quotient rule right?

OpenStudy (accessdenied):

yeah, you could use quotient rule

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

can someone get me started please

OpenStudy (anonymous):

dont want the answer. just get me started.

OpenStudy (accessdenied):

\[ F_1 = s + 4, F_2 = s^2 + s + 14 \] \[ \frac{F_1' F_2 - F_1 F_2 '}{F_2^2} \]

OpenStudy (anonymous):

do u think all of these rules will be given to me on a test or should i memorize them?

OpenStudy (accessdenied):

i doubt they'd be given, it'd be better to memorize them. usually (at least when i was learning the rules), they start to become more second nature as you do more problems.

OpenStudy (anonymous):

true. too much to store in the brain though. haha

OpenStudy (accessdenied):

its technically possible to use product rule on those quotients of functions if you redefine it in the form f/g = (f)(g)^(-1) , that could possibly help in minimizing the formulas you need to remember... lol

OpenStudy (anonymous):

ok im getting the wrong final answer for this. can you help me out here?

OpenStudy (accessdenied):

what are you getting for your answer?

OpenStudy (anonymous):

-s^2 - 8s + 10 / (s^2 + s + 14)^2

OpenStudy (anonymous):

maybe im making a syntax error

OpenStudy (accessdenied):

I'm gonna use f(s) = s+4 and g(s) = s^2 + s + 14 instead of my previous F1 and F2 for easier reading. Given those f(s) and g(s), their derivatives... f'(s) = 1 g'(s) = 2s + 1 f'(s)g(s) - f(s)g'(s) -------------- gIs)^2 maybe that'd be easier to understand

OpenStudy (accessdenied):

oh, hmm, i get that too

OpenStudy (accessdenied):

maybe it would like you to multiply out the denominator

OpenStudy (anonymous):

i made a syntax error. its an old school hw website.

OpenStudy (accessdenied):

ah, i see